Mathematics 242 – Principles of Analysis Definitions Quiz 3 – September 23, 2005

Your Name:

1. What does it mean for a set to be countable?

2. What is the precise definition of the statement “the sequence (x_{n}) converges to the real number a” (or equivalently, lim_{n→∞} x_{n} = a)?